. An airplane is flying at 800 km/h at an altitude where the

pressure is 46.57 kPa and the temperature is -24.6°C. The airplane is powered

by a jet engine having a maximum cycle pressure of 280 kPa and a maximum

cycle temperature of 1090°C. The air flow is 95 kg/s. Determine the maximum

propulsive force that can be produced by the engine.


1
Expert's answer
2021-07-20T15:19:06-0400

We know that

"F=\frac{dP}{dt}+v\frac{dm}{dt}"

Put value


"F=({280-46.57})\times10^3+\frac{800\times10^3}{3600}\times95=233430+21111.11=254541.11N"

F=254.54KN